# Compositional Zero-shot Learning via Progressive Language-based Observations

### v1

@@ 3.2 PLO-VLM
As illustrated in Figure 2(a), PLO-VLM adopts a dynamic two-step observation framework to decide whether to prioritize observing the state or object primitive. Next, the prompt features of the selected primitive are integrated into the image using cross-modal attention (CA), which highlights regions of interest related to the observation. By further comparing the similarities between the refined image features and the text features of the remaining primitives, PLO-VLM effectively recognizes state-object compositions.
@@
1) Pre-observing. The pre-observing classifier recognizes the first primitive by measuring the cosine similarity $\mathcal{S}(\cdot,\cdot)$ between the image representation $\mathbf{v}$ and the text representations of the state primitive $\mathbf{t}_{s}$ and the object primitive $\mathbf{t}_{o}$:
$\mathcal{F}^{pre}_{obs}(\mathbf{v},\mathbf{t}_{s},\mathbf{t}_{o})=\mathcal{S}(\mathbf{v},\mathbf{t}_{s})\oplus\mathcal{S}(\mathbf{v},\mathbf{t}_{o}),$ (2)
@@
where $\oplus$ denotes concatenation. The first observation is the prompt of $argmax(\mathcal{F}^{pre}_{obs}(\cdot))$. Its corresponding text representation is $\mathbf{t}^{pre}_{obs}$. For instance in Figure 2(a), the first observation is the prompt of the state “mashed object”, and $\mathbf{t}^{pre}_{obs}$ will correspond to the text representation $\mathbf{t}_{s}$.
@@
2) Post-observing. After obtaining the first observation, we employ learnable residual cross-modal attention modules to extract the interest features of the first observation for refined image representations $\widetilde{\mathbf{v}}$, which are widely adopted in existing methods [3, 10, 22]. The CA module is defined as:
$\text{CA}(\mathbf{q},\mathbf{K},\mathbf{V})=\mathbf{q}+\rm{FFN}(LN(\mathbf{q}+MHA(\mathbf{q},\mathbf{K},\mathbf{V}))),$ (3)
@@
where $\mathbf{q}$, $\mathbf{K}$, and $\mathbf{V}$ are the features of query, key, and value, respectively. The module FFN, LN, and MHA are feed-forward networks, layer normalization, and multi-head attention, respectively. The refined image representations $\widetilde{\mathbf{v}}$ are derived based on the first observation as follows:
$\widetilde{\mathcal{V}}=\text{CA}(\mathcal{V},\mathcal{T}^{pre}_{obs},\mathcal{T}^{pre}_{obs}),\widetilde{\mathbf{v}}=\text{FC}(\widetilde{\mathcal{V}}),$ (4)
@@
$p(s|I)$ $=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs}),\tau),\text{ if }\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{s},$ (5)
@@
where $\pi(\cdot)$ is the softmax function and $\tau$ denotes the temperature hyper-parameter. The $\mathbf{t}^{post}_{obs}$ represents the text presentation of the prompt of retaining state or object primitive, respectively. Additionally, to ensure that the refined feature encompasses both state and object features, we also compare its similarity with the text presentations of the composition prompt, which is defined as follows:
$p(c|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}_{c}),\tau),$ (6)
@@
3.3 PLO-LLM
As shown in Figure 2(b), it employs a multi-step observation process and starts by generating a sequence of observation prompts for each composition category using an LLM, i.e., GPT [4]. Then, the final compositional classification is determined by observing step by step whether observation prompts exist in the image.
1) Observation Prompts Generation. To elaborate, for each composition category $c\in C$, a sequence of the observation hard prompts $\mathcal{P}_{c}$ = $\{\mathbf{P}_{c}^{(i)}\}^{n}_{i=1}$ is generated using an LLM with the designed prompt (refer to the Appendix), where $n$ denotes the number of observation steps. Note that these prompts are frozen. For instance, in Figure 2(b), the observation prompts are listed in the blue box, e.g., “yellow, mushy substance” for “mashed banana”.
2) Step-by-step Observing. At each observation step $i$, the image representation $\mathbf{v}^{(i)}$ and the text representation of the current observation prompt $\mathbf{t}^{(i)}_{obs}$ are fed into the learnable CA module. The CA module updates the image representation to produce the refined image representation $\widetilde{\mathbf{v}}^{(i)}$:
$\widetilde{\mathcal{V}}^{(i)}$ $=\text{CA}(\mathcal{V}^{(i)},\mathcal{T}^{(i)}_{obs},\mathcal{T}^{(i)}_{obs}),$ (7)
$\mathbf{v}^{(i+1)}$ $=\widetilde{\mathbf{v}}^{(i)}=\text{FC}(\widetilde{\mathcal{V}}^{(i)}),$
where $\widetilde{\mathbf{v}}^{(i)}$ incorporates information from the current observation prompt, which is then utilized in the next observation step to calculate the similarity between the image and the next observation prompt $\mathbf{t}^{(i+1)}_{obs}$. This process is repeated for all observation steps, iteratively updating the image representation with information from each observation prompt. The result of each observation is $\mathcal{S}(\mathbf{v}^{(i)},\mathbf{t}^{(i)}_{obs})$.
@@ 3.3 PLO-LLM
$p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)})=\pi(\mathcal{S}(\mathbf{v}^{(i)},\mathbf{t}^{(i)}_{obs}),\tau).$ (8)
Furthermore, to facilitate the model in learning semantic representations suitable for CZSL without being solely influenced by hard prompts, we also incorporate the similarity between the original image representation $\mathbf{v}$ and the text representation $\mathbf{t}_{c}$ of the learnable composition soft prompts to aid in prediction. The probability is calculated as follows:
$p^{Soft}(c|I)=\pi(\mathcal{S}(\mathbf{v},\mathbf{t}_{c}),\tau).$ (9)
3.4 Training Objectives and Inference
$\mathcal{L}_{PLO}^{VLM}=\mathcal{L}_{obs}+\mathcal{L}_{s}+\mathcal{L}_{o}+\mathcal{L}_{c}.$ (10)
$\mathcal{L}_{obs}=$ $-y_{obs}\log(\sigma(\mathcal{F}^{pre}_{obs}))$ (11)
$+(1-y_{obs})\log(1-\sigma(\mathcal{F}^{pre}_{obs})),$
where $\sigma(\cdot)$ denotes the sigmoid function. ${y}_{obs}$ = $[y_{s},y_{o}]$ is the ground-truth multi-label target for the pre-observing classifier, where $y_{s}$ and $y_{o}$ are binary values of ground-truth state and object categories, respectively.
The optimization of classification involves three independent cross-entropy losses: $\mathcal{L}_{s}$ for state, $\mathcal{L}_{o}$ for object, and $\mathcal{L}_{c}$ for composition. Taking the state classification as an illustrative example, its loss $\mathcal{L}_{s}$ is defined as:
$\mathcal{L}_{s}=-\textstyle{\sum}_{S}\;y_{s}log(p(s|I)).$ (12)
During inference, we pick the composition with the highest score of $p(c|I)$ as our predicted label for the input image.
PLO-LLM. In the training of PLO-LLM, the loss includes the cross-entropy loss $\mathcal{L}_{step}$ for each observation step and cross-entropy loss $\mathcal{L}_{c}$ for composition:
$\mathcal{L}_{PLO}^{LLM}=\mathcal{L}_{step}+\mathcal{L}_{c},$ (13)
$\!\!\mathcal{L}_{step}\!=\!-\!\sum_{i=1}^{n}\sum_{j\in C}{y^{(i)}_{c}}\log(p(c_{j}^{(i)}|I,c_{j}^{(1)},\dots,c_{j}^{(i-1)})),$ (14)
where ${y}^{(i)}_{c}$ is the ground truth one-hot encoded composition label for the $i$-th observation prompt at step $i$.
Since PLO-LLM includes both the step-by-step hard prompt observation and soft prompt during training, the inference process involves predicting the composition label $\hat{y}_{c}$ by calculating the probabilities of the composition categories using the following equations:
$p(c|I)$ $=p^{Soft}(c|I)+p^{Hard}(c|I),$ (15)
$p^{Hard}(c|I)$ $=\prod_{i=1}^{n}p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)}),$
where $p(c|I)$ is the probability of composition category $c$, $p^{Soft}(c|I)$ is the probability based on soft prompts, and $p^{Hard}(c|I)$ is based on hard prompts at each observing step. The composition category with the highest overall probability is chosen as the final predicted composition label $\hat{y}_{c}$.
@@ 3.4 Training Objectives and Inference
Closed-World CLIP [34] ${}_{\text{ICML'21}}$ 30.2 46.0 26.1 11.0 15.8 49.1 15.6 5.0 7.5 25.0 8.6 1.4
CoOp [43] ${}_{\text{IJCV'22}}$ 34.4 47.6 29.8 13.5 52.1 49.3 34.6 18.8 20.5 26.8 17.1 4.4
Co-CGE [26] ${}_{\text{TPAMI'22}}$ 46.7 45.9 33.1 17.0 63.4 71.3 49.7 36.3 34.1 21.2 18.9 5.7
ProDA [24] ${}_{\text{CVPR'22}}$ 37.4 51.7 32.7 16.1 63.7 60.7 47.6 32.7 - - - -
PromptCompVL [24] ${}_{\text{arXiv'22}}$ 48.5 47.2 35.3 18.3 64.4 64.0 46.1 32.2 - - - -
CSP [30] ${}_{\text{ICLR'23}}$ 46.6 49.9 36.3 19.4 64.2 66.2 46.6 33.0 28.8 26.8 20.5 6.2
DFSP(i2t) [22] ${}_{\text{CVPR'23}}$ 47.4 52.4 37.2 20.7 64.2 66.4 45.1 32.1 35.6 29.3 24.3 8.7
DFSP(BiF) [22] ${}_{\text{CVPR'23}}$ 47.1 52.8 37.7 20.8 63.3 69.2 47.1 33.5 36.5 32.0 26.2 9.9
DFSP(t2i) [22] ${}_{\text{CVPR'23}}$ 46.9 52.0 37.3 20.6 66.7 71.7 47.2 36.0 38.2 32.0 27.1 10.5
Troika [10] ${}_{\text{arXiv'23}}$ 49.0 53.0 39.3 22.1 66.8 73.8 54.6 41.7 41.0 35.7 29.4 12.4
PLID [3] ${}_{\text{arXiv'23}}$ 49.7 52.4 39.0 22.1 67.3 68.8 52.4 38.7 38.8 33.0 27.9 11.0
GIPCOL [38] ${}_{\text{WACV'24}}$ 48.5 49.6 36.6 19.9 65.0 68.5 48.8 36.2 31.9 28.4 22.5 7.1
PLO-VLM (Ours) 49.6 52.7 39.0 22.2 67.8 75.6 53.1 42.0 43.9 38.2 32.2 14.5
@@
Open-World CLIP [34] ${}_{\text{ICML'21}}$ 30.1 14.3 12.8 3.0 15.7 20.6 11.2 2.2 7.5 4.6 4.0 0.3
CoOp [43] ${}_{\text{IJCV'22}}$ 34.6 9.3 12.3 2.8 52.1 31.5 28.9 13.2 21.0 4.6 5.5 0.7
Co-CGE [26] ${}_{\text{TPAMI'22}}$ 38.1 20.0 17.7 5.6 59.9 56.2 45.3 28.4 33.2 3.9 5.3 0.9
ProDA [24] ${}_{\text{CVPR'22}}$ 37.5 18.3 17.3 5.1 63.9 34.6 34.3 18.4 - - - -
PromptCompVL [24] ${}_{\text{arXiv'22}}$ 48.5 16.0 17.7 6.1 64.6 44.0 37.1 21.6 - - - -
CSP [30] ${}_{\text{ICLR'23}}$ 46.3 15.7 17.4 5.7 64.1 44.1 38.9 22.7 28.7 5.2 6.9 1.2
DFSP(i2t) [22] ${}_{\text{CVPR'23}}$ 47.2 18.2 19.1 6.7 64.3 53.8 41.2 26.4 35.6 6.5 9.0 2.0
DFSP(BiF) [22] ${}_{\text{CVPR'23}}$ 47.1 18.1 19.2 6.7 63.5 57.2 42.7 27.6 36.4 7.6 10.6 2.4
DFSP(t2i) [22] ${}_{\text{CVPR'23}}$ 47.5 18.5 19.3 6.8 66.8 60.0 44.0 30.3 38.3 7.2 10.4 2.4
Troika [10] ${}_{\text{arXiv'23}}$ 48.8 18.7 20.1 7.2 66.4 61.2 47.8 33.0 40.8 7.9 10.9 2.7
PLID [3] ${}_{\text{arXiv'23}}$ 49.1 18.7 20.0 7.3 67.6 55.5 46.6 30.8 39.1 7.5 10.6 2.5
GIPCOL [38] ${}_{\text{WACV'24}}$ 48.5 16.0 17.9 6.3 65.0 45.0 40.1 23.5 31.6 5.5 7.3 1.3
PLO-VLM (Ours) 49.5 18.7 20.5 7.4 68.0 63.5 47.8 33.1 43.9 10.4 13.9 3.9
Table 1: Performance (%) of state-of-the-art CZSL models on three datasets in closed-world and open-world settings.
The training and inference procedures of PLO-VLM and PLO-LLM are shown in the Appendix.
4 Experiments
4.1 Experiment Settings

### v2

@@ 3.2. PLO-VLM
As shown in Figure 2(a), PLO-VLM adopts a dynamic two-step observation framework to decide whether to prioritize observing the state or object primitive concept. Next, the prompt features of the selected primitive are integrated into the image using cross-modal attention (CA), which highlights regions of interest related to the observation. By further comparing the similarities between the refined image features and the text features of the remaining primitives, PLO-VLM effectively recognizes state-object compositions.
@@
1) Pre-observing. The pre-observing classifier recognizes the first primitive by measuring the cosine similarity $\mathcal{S}(\cdot,\cdot)$ between the image representation $\mathbf{v}$ and the text representations of the state primitive $\mathbf{t}_{s}$ and object primitive $\mathbf{t}_{o}$:
(2) $\mathcal{F}^{pre}_{obs}(\mathbf{v},\mathbf{t}_{s},\mathbf{t}_{o})=\mathcal{S}(\mathbf{v},\mathbf{t}_{s})\oplus\mathcal{S}(\mathbf{v},\mathbf{t}_{o}),$
@@
where $\oplus$ denotes concatenation. The first observation is the prompt of $argmax(\mathcal{F}^{pre}_{obs}(\cdot))$. Its corresponding text representation is $\mathbf{t}^{pre}_{obs}$. For instance in Figure 2(a), the first observation is the prompt of the state “mashed object”, and $\mathbf{t}^{pre}_{obs}$ is text representation $\mathbf{t}_{s}$.
@@
2) Post-observing. After obtaining the first observation, we use learnable residual CA modules to extract interest features of the first observation for refined image representations $\widetilde{\mathbf{v}}$, which are widely used in methods (Huang et al., 2023; Lu et al., 2023a). The CA module is defined as:
(3) $\text{CA}(\mathbf{q},\mathbf{K},\mathbf{V})=\mathbf{q}+\rm{FFN}(LN(\mathbf{q}+MHA(\mathbf{q},\mathbf{K},\mathbf{V}))),$
@@
where $\mathbf{q}$, $\mathbf{K}$, and $\mathbf{V}$ are the features of query, key, and value. The module FFN, LN, and MHA are feed-forward networks, layer normalization, and multi-head attention. The refined image representations $\widetilde{\mathbf{v}}$ are derived based on the first observation as:
[…]
@@
(5) $p(s|I)$ $=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}^{post}_{obs}),\tau),\text{ if }\mathbf{t}^{post}_{obs}\leftarrow\mathbf{t}_{s},$
@@
where $\pi(\cdot)$ is the softmax function and $\tau$ denotes the temperature hyper-parameter. The $\mathbf{t}^{post}_{obs}$ represents the text presentation of the prompt of retaining state or object primitive, respectively. Besides, to ensure that the refined feature encompasses both state and object features, we also compare its similarity with the text presentations of the composition prompt, which is written as:
(6) $p(c|I)=\pi(\mathcal{S}(\widetilde{\mathbf{v}},\mathbf{t}_{c}),\tau),$
@@
3.3. PLO-LLM
In Figure 2(b), we employ a multi-step observation process and starts by directly generating a sequence of observation cues (graduated descriptions) for each composition category using an LLM, i.e., GPT (Brown et al., 2020). Then, the final compositional classification is determined by observing step by step whether cues exist in the image.
1) Observation Cues Generation. To elaborate, for each composition category $_{\!}c\in_{\!}C$, a sequence of the hard prompts (observation cues) $\mathcal{P}_{c}$ = $\{\mathbf{P}_{c}^{(i)}\}^{n}_{i=1}$ is generated using an LLM with the designed prompt (refer to the Appendix), where $n$ denotes the number of observation steps. Note that these prompts are frozen. For instance, in Figure 2(b), the observation cues are listed in the blue box, e.g., “yellow, mushy substance” for “mashed banana”.
2) Step-by-step Observing. At each observation step $i$, the image representation $\mathbf{v}^{(i)}$ and the text representation of the current observation cue $\mathbf{t}^{(i)}_{obs}$ are fed into the learnable CA module. The CA module updates the image representation to produce the refined image representation $\widetilde{\mathbf{v}}^{(i)}$:
(7) $\widetilde{\mathcal{V}}^{(i)}\!=\!\text{CA}(\mathcal{V}^{(i)},\mathcal{T}^{(i)}_{obs},\mathcal{T}^{(i)}_{obs}),\quad\mathbf{v}^{(i+1)}=\widetilde{\mathbf{v}}^{(i)}=\text{FC}(\widetilde{\mathcal{V}}^{(i)}),$
where $\widetilde{\mathbf{v}}^{(i)}$ incorporates information from the current observation cue, which is then utilized in the next observation step to calculate the similarity between the image and the next observation cue $\mathbf{t}^{(i+1)}_{obs}$. This process is repeated for all observation steps, iteratively updating the image representation with information from each observation cue. The result of each observation is $\mathcal{S}(\mathbf{v}^{(i)},\mathbf{t}^{(i)}_{obs})$.
@@ 3.3. PLO-LLM
(8) $p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)})=\pi(\mathcal{S}(\mathbf{v}^{(i)},\mathbf{t}^{(i)}_{obs}),\tau).$
Furthermore, to facilitate the model in learning semantic representations suitable for CZSL without being solely influenced by hard prompts, we also incorporate the similarity between the original image representation $\mathbf{v}$ and the text representation $\mathbf{t}_{c}$ of the learnable composition soft prompts. The probability is calculated as:
(9) $p^{Soft}(c|I)=\pi(\mathcal{S}(\mathbf{v},\mathbf{t}_{c}),\tau).$
@@
Closed-World CLIP (Radford et al., 2021) $_{\text{ICML'21}}$ 30.2 46.0 26.1 11.0 15.8 49.1 15.6 5.0 7.5 25.0 8.6 1.4
CoOp (Zhou et al., 2022b) $_{\text{IJCV'22}}$ 34.4 47.6 29.8 13.5 52.1 49.3 34.6 18.8 20.5 26.8 17.1 4.4
Co-CGE (Mancini et al., 2022) $_{\text{TPAMI'22}}$ 46.7 45.9 33.1 17.0 63.4 71.3 49.7 36.3 34.1 21.2 18.9 5.7
ProDA (Lu et al., 2022) $_{\text{CVPR'22}}$ 37.4 51.7 32.7 16.1 63.7 60.7 47.6 32.7 - - - -
CSP (Nayak et al., 2023) $_{\text{ICLR'23}}$ 46.6 49.9 36.3 19.4 64.2 66.2 46.6 33.0 28.8 26.8 20.5 6.2
DFSP(i2t) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 47.4 52.4 37.2 20.7 64.2 66.4 45.1 32.1 35.6 29.3 24.3 8.7
DFSP(BiF) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 47.1 52.8 37.7 20.8 63.3 69.2 47.1 33.5 36.5 32.0 26.2 9.9
DFSP(t2i) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 46.9 52.0 37.3 20.6 66.7 71.7 47.2 36.0 38.2 32.0 27.1 10.5
GIPCOL (Xu et al., 2023) $_{\text{WACV'24}}$ 48.5 49.6 36.6 19.9 65.0 68.5 48.8 36.2 31.9 28.4 22.5 7.1
CAILA (Li et al., 2024) $_{\text{WACV'24}}$ 51.4 53.3 39.7 23.2 66.8 72.5 56.0 42.5 44.3 35.8 31.4 13.9
Troika (Huang et al., 2023) $_{\text{CVPR'24}}$ 49.0 53.0 39.3 22.1 66.8 73.8 54.6 41.7 41.0 35.7 29.4 12.4
CDS-CZSL (Li et al., 2024) $_{\text{CVPR'24}}$ 50.3 52.9 39.2 22.4 63.9 74.8 52.7 39.5 38.3 34.2 28.1 11.1
PLID (Bao et al., 2024) $_{\text{ECCV'24}}$ 49.7 52.4 39.0 22.1 67.3 68.8 52.4 38.7 38.8 33.0 27.9 11.0
PLO-VLM (Ours) 49.7 52.8 39.4 22.3 67.8 75.6 53.1 42.0 43.9 38.2 32.2 14.5
@@
PLO-VLM (Ours) 51.6 53.7 40.2 23.4 70.3 75.8 55.3 43.6 44.7 38.1 33.0 14.9
Open-World CLIP (Radford et al., 2021) $_{\text{ICML'21}}$ 30.1 14.3 12.8 3.0 15.7 20.6 11.2 2.2 7.5 4.6 4.0 0.3
CoOp (Zhou et al., 2022b) $_{\text{IJCV'22}}$ 34.6 9.3 12.3 2.8 52.1 31.5 28.9 13.2 21.0 4.6 5.5 0.7
Co-CGE (Mancini et al., 2022) $_{\text{TPAMI'22}}$ 38.1 20.0 17.7 5.6 59.9 56.2 45.3 28.4 33.2 3.9 5.3 0.9
ProDA (Lu et al., 2022) $_{\text{CVPR'22}}$ 37.5 18.3 17.3 5.1 63.9 34.6 34.3 18.4 - - - -
CSP (Nayak et al., 2023) $_{\text{ICLR'23}}$ 46.3 15.7 17.4 5.7 64.1 44.1 38.9 22.7 28.7 5.2 6.9 1.2
DFSP(i2t) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 47.2 18.2 19.1 6.7 64.3 53.8 41.2 26.4 35.6 6.5 9.0 2.0
DFSP(BiF) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 47.1 18.1 19.2 6.7 63.5 57.2 42.7 27.6 36.4 7.6 10.6 2.4
DFSP(t2i) (Lu et al., 2023a) $_{\text{CVPR'23}}$ 47.5 18.5 19.3 6.8 66.8 60.0 44.0 30.3 38.3 7.2 10.4 2.4
GIPCOL (Xu et al., 2023) $_{\text{WACV'24}}$ 48.5 16.0 17.9 6.3 65.0 45.0 40.1 23.5 31.6 5.5 7.3 1.3
CAILA (Li et al., 2024) $_{\text{WACV'24}}$ 51.4 20.1 20.9 8.0 65.1 59.6 44.8 29.9 43.8 11.4 7.9 3.3
Troika (Huang et al., 2023) $_{\text{CVPR'24}}$ 48.8 18.7 20.1 7.2 66.4 61.2 47.8 33.0 40.8 7.9 10.9 2.7
CDS-CZSL (Li et al., 2024) $\color[rgb]{.5,.5,.5}{}_{\text{CVPR'24}}$} 49.4 21.8 22.1 8.5 64.7 61.3 48.2 32.3 37.6 8.2 11.6 2.7
PLID (Bao et al., 2024) $_{\text{ECCV'24}}$ 49.1 18.7 20.0 7.3 67.6 55.5 46.6 30.8 39.1 7.5 10.6 2.5
PLO-VLM (Ours) 49.7 19.4 21.4 7.8 68.0 63.5 47.8 33.1 43.9 10.4 13.9 3.9
PLO-LLM (Ours) 49.8 19.1 20.6 7.5 68.0 58.2 48.8 33.8 43.6 10.4 13.6 3.9
Table 1. Performance comparison (%) on three datasets. $\dagger$ denotes uses MoA adapter (Zheng et al., 2024). $\star$ indicates the use of an additional filter trick w/o releasing codes. $\ast$ denotes reproduction results using the official released code.
3.4. Training Objectives and Inference
(10) $\mathcal{L}_{PLO}^{VLM}=\mathcal{L}_{obs}+\mathcal{L}_{s}+\mathcal{L}_{o}+\mathcal{L}_{c}.$
(11) $\mathcal{L}_{obs}=-y_{obs}\log(\sigma(\mathcal{F}^{pre}_{obs}))+(1-y_{obs})\log(1-\sigma(\mathcal{F}^{pre}_{obs})),$
where $\sigma(\cdot)$ is sigmoid function. ${y}_{obs}$ = $y_{s},y_{o}]$ is ground-truth multi-label target for the pre-observing classifier, where $y_{s}$ and $y_{o}$ are binary values of ground-truth state and object categories. Both $\mathcal{F}^{pre}_{obs}$ and ${y}_{obs}$ with the shape of [—SS— + — OO—], where —SS— and —OO— are the total class numbers of states and objects.
The optimization of classification involves three independent cross-entropy losses: $\mathcal{L}_{s}$ for state, $\mathcal{L}_{o}$ for object, and $\mathcal{L}_{c}$ for composition. Taking the state as example, its loss $\mathcal{L}_{s}$ is defined as:
(12) $\mathcal{L}_{s}=-\textstyle{\sum}_{S}\;y_{s}log(p(s|I)).$
During inference, we pick the composition with the highest score of $p(c|I)$ as our predicted label.
PLO-LLM. The loss includes the cross-entropy loss $\mathcal{L}_{step}$ for each observation step and cross-entropy loss $\mathcal{L}_{c}$ for composition:
(13) $\mathcal{L}_{PLO}^{LLM}=\mathcal{L}_{step}+\mathcal{L}_{c}.$
(14) $\mathcal{L}_{step}=-\sum_{i=1}^{n}\sum_{j\in C}{y^{(i)}_{c}}\log(p(c_{j}^{(i)}|I,c_{j}^{(1)},\dots,c_{j}^{(i-1)})),$
where ${y}^{(i)}_{c}$ is the ground truth one-hot encoded composition label for the $i$-th observation at step $i$.
Since PLO-LLM includes both the step-by-step hard prompt observation and soft prompt during training, the inference involves predicting the composition label $\hat{y}_{c}$ by calculating the probabilities of the composition categories using the following equations:
(15) $p(c|I)=p^{Soft}(c|I)+p^{Hard}(c|I),$
$p^{Hard}(c|I)=\prod_{i=1}^{n}p(c^{(i)}|I,c^{(1)},\dots,c^{(i-1)}),$
where $p(c|I)$ is the probability of the composition category $c$ and $p^{Soft}(c|I)$ is the probability based on soft prompts, and $p^{Hard}(c|I)$ is based on hard prompts at each observing step. The composition category with the highest probability is chosen as the final predicted composition label $\hat{y}_{c}$. The training and inference procedures of PLO-VLM and PLO-LLM are shown in the Appendix.
4. Experiments
4.1. Experiment Settings